OpenStudy (anonymous):

Q26. If f(x) = sin(x) for all x, then the average value of f on the interval [0, π] is (A) ½ (B) 1/π (C) π/2 (D) 2/π

OpenStudy (amistre64):

hmm.... 0+1+0 ------ = 1/3 3 0+.5+1+.5+0 ------------- =2/5 5 are you sure the endpoints are included?

OpenStudy (anonymous):

ya